# How Do Social Bots Participate in Misinformation Spread? A Comprehensive Dataset and Analysis

### v2

@@ 4 Misinformation and Social Bots
Social bots are deeply involved in information spread
We check the bot percentage in MisBot:
@@
Misinformation with the same topics has similar content, providing the basis of echo chambers, and social bots amplify this phenomenon
@@
We first group all pieces of misinformation into clusters with the same topics according to the judgment [^13] We provide the clustering algorithm in Appendix C.1.. We group 11,393 pieces of misinformation into 2,270 clusters, each of which represents a specific topic or event, e.g., “The last two minutes of the air crash”. We aim to explore the textual content similarity of misinformation with the same topics and across different topics.
@@
We first visualize the misinformation content representations in Figure 4, where we select 10 topics with the highest number of misinformation items. It shows the BERT representation of misinformation content using t-SNE dimensionality reduction [^14] We provide a detailed description of each selected topic in Appendix C.2.. The silhouette score is 0.29, illustrating that the clusters are cohesive. It indicates that each cluster shares similar content while different clusters share significant differences. It initially suggests that the misinformation environment is homogeneous, providing the basis for echo chambers.
@@
Figure 4: Visualization of misinformation content representations within the top 10 topics. Each dot corresponds to a piece of misinformation, colored according to its topic. The topic labels annotated by the judgment are plotted at each cluster center. We also calculate the silhouette score (×$\times 100$). The cohesive clusters indicate misinformation about the same topic having similar content, providing the basis of echo chambers.
@@
We conduct further numerical analysis by calculating semantic-level and token-level pairwise scores between any two clusters, where higher scores mean the content of the two clusters is more similar. For semantic level, we employ the cosine similarity of the BERT representations, while for token level, we leverage the ROUGE-L score [^15] We provide the experiment settings and detailed analysis in Appendix C.3.. Figure 16 in Appendix C.4 presents the heatmap of the pairwise scores. It illustrates that the values on the diagonal are much greater. For semantic level, the average diagonal value is 0.9448, and the others’ average is 0.5847. For token level, the average diagonal value is 0.781,5 and the others’ average is 0.0773. The numerical results emphasize that misinformation with the same topics has similar content, and misinformation with different topics has distinct content.
@@
Then we explore how social bots participate in misinformation spread. Inspired by existing works, we consider an account a potential echo chamber member if it engages in at least two misinformation spreads in the same cluster. Figure 5 displays the distribution of bot percentage among echo chamber members and non-members within various clusters. It illustrates that around 18% non-members are social bots. Meanwhile, the members do not contain bots in about half of the clusters. However, in the other half, members exhibit a higher bot percentage across most clusters compared to non-members, reaching up to 50% in many clusters. We speculate that social bots might reinforce the echo chamber effect by actively engaging in misinformation spread with the same topics.
@@
Social bots generate similar content aiming to manipulate public opinions
Online information consumers are reluctant to process information deliberately (Möller et al., 2020), becoming susceptible to cognitive biases (Pennycook et al., 2018; Vosoughi et al., 2018). We aim to explore how public opinion changes and how social bots potentially manipulate it. We focus on how public sentiments and stances change in MisBot [^16] We provide information on obtaining the sentiments and stances of a social text in Appendix C.5..
  • For public sentiments, we categorize sentiments into neutral and non-neutral (including happy, angry, surprised, sad, and fearful). Figure 17 in Appendix C.6 presents sentiment distribution in different social texts. It illustrates that misinformation would publish more emotional content, while real information would naturally be reported. On the other hand, public reactions are always emotional, where misinformation shows more anger while real information shows more happiness. Thus, public sentiments are emotional. We further explore the degree or extent to which public sentiments change over the information spread, introducing a variation measure [^17] We provide the detailed introduction of this measure in Appendix C.7.:
@@
    where $f(x_{k})$ denotes neutral sentiment proportion at time $x_{k}$. Figure 6 visualizes sentiment variation distribution, where a larger value means a more drastic change. The average values of misinformation and real information are 0.287 and 0.225. It illustrates that public sentiment changes are dramatic during information spread.
@@
    Figure 7: The proportion of comments with different stances as the comments increase. Public stances become increasingly polarized, where misinformation contains more comments with clear stances.
@@
Therefore, we can conclude that as the information spreads, public opinions, including sentiment and stance, become polarized, especially regarding misinformation. We then numerically prove the correlation between polarization and social bots by the Pearson correlation coefficient [^18] We provide the calculation method and more analysis in Appendix C.8.: the number of social bots demonstrates strong correlations with the number of comments with non-neutral stances (r = 0.6661) and sentiments (r = 0.6750). The relatively high correlation coefficients indicate that social bots might influence public opinion.
Figure 8: Distribution of social bots’ and humans’ semantic similarities, where social bots present higher similarities. Namely, social bots would publish similar content to manipulate public opinions.
We further explore social bot characteristics in information spread. Figure 8 presents the semantic similarity of social bots and genuine accounts [^19] We provide the calculation method in Appendix C.9.. It illustrates that social bots present higher values. Namely, social bots would publicize similar content and amplify the bandwagon effect, where online users adopt behaviors or actions simply because others are doing so, influencing the information spread (Wang and Zhu, 2019). Figure 18 in Appendix C.10 illustrates the stances and sentiments of comments published by social bots. It demonstrates that social bots publish more emotional content and comments with clear stances. The results enhance the finding that social bots generate similar content, aiming to manipulate public opinions.
@@ 5.1 Misinformation Detection
Mainstream detectors focuses on the information content, including text (Hartl and Kruschwitz, 2022; Xiao et al., 2024; Gong et al., 2024), images (Liu et al., 2023a; Zhang et al., 2024b, d), and videos (Tan et al., 2023a; Bu et al., 2024). They extract features such as emotion (Zhang et al., 2021) and employ neural networks such as graph neural networks (Tao et al., 2024; Zhang et al., 2024f; Lu et al., 2024) or neurosymbolic reasoning (Dong et al., 2024) to characterize information. Besides information content, the context, such as user interactions (Shu et al., 2019; Lu and Li, 2020), user profile (Sun et al., 2023; Xu et al., 2024), and evidence (Chen et al., 2024) provide helpful signals to detect misinformation. These models would model propagation patterns (Cui and Jia, 2024), construct news environments (Yin et al., 2024), or extract multi-hop facts (Zhang et al., 2024a) to enhance detection performance. Recently, to combat LLM-generated misinformation (Zhang et al., 2024e; Venkatraman et al., 2024), models employing LLMs (Wan et al., 2024; Nan et al., 2024) through prompting (Guan et al., 2024; Hu et al., 2024) and in-context learning (Wang et al., 2024) have been proposed.
@@ 6 Conclusion
We explored the interplay between misinformation and social bots. We proposed a novel dataset named MisBot containing information and annotations of misinformation and social bots. MisBot is the most comprehensive, and misinformation and real information are distinguishable, and social bots have high annotation quality. The analysis illustrates that (i) social bots are deeply involved in information spread; (ii) misinformation provides the basis of echo chambers, and social bots amplify this phenomenon; (iii) social bots generate similar content aiming to manipulate public opinions.
@@
Our proposed dataset is the largest containing misinformation and social bot annotations simultaneously. Meanwhile, it contains multiple modalities of news, including images and video,s and interactions of users, including comments and reposts. However, due to the focus on news spread, it does not contain interactions like the friend relationship, missing potential relations between social bots and genuine accounts. Meanwhile, we propose a weak-supervised framework to annotate social bots, whose accuracy is similar to crowd-sourcing. However, it struggles to achieve better recall and might miss several social bots. Finally, the experiments in this work focus primarily on the Sina Weibo platform. We expect to expand our experiments and analysis to other social media platforms such as $\mathbb{X}$ (Twitter) or Reddit, in future work.
@@
The research of misinformation and social bots is essential in countering online malicious content. This research demonstrates that social bots would amplify the spread of misinformation, enhancing echo chambers and manipulating public opinions. However, this work may increase the risk of dual-use, where malicious actors may develop advanced social bots to spread misinformation. We will establish controlled access to ensure that the data and trained model checkpoint are only publicly available to researchers. Meanwhile, we will hide the privacy information in the dataset when we publish it.
@@ A.4 Entity Filter
A.5 Verified Accounts
@@ A.8 Annotation Guideline
Based on the criteria, we write a guideline document for human annotators in Figures 12 and 13. Each human annotator must read this document before annotating.
@@ B.3 Misinformation Detector
where $\mathrm{mean}(\cdot)$ denotes the meaning operator. To encode video, we sample 256 frames from each video and resize each frame into $224\times 224$. We employ pre-trained VideoMAE [^28] Here is the model link. (Tong et al., 2022) $\mathrm{VideoMAE}(\cdot)$ For each time step, we take 16 frames and set the interval to 12 frames. We could obtain:
@@ B.4 Detector Ablation Study
w/o Comment $\mathop{93.0^{\star}_{\pm 1.4}}\limits_{2.3\%\downarrow}$ $\mathop{89.5^{\star}_{\pm 1.6}}\limits_{3.0\%\downarrow}$ $\mathop{86.1^{\star}_{\pm 3.9}}\limits_{8.1\%\downarrow}$ $\mathop{93.3^{\dagger}_{\pm 1.6}}\limits_{2.6\%\uparrow}$
w/o Repost $\mathop{89.4^{\star}_{\pm 2.0}}\limits_{6.0\%\downarrow}$ $\mathop{85.3^{\star}_{\pm 2.2}}\limits_{7.6\%\downarrow}$ $\mathop{76.8^{\star}_{\pm 4.0}}\limits_{18.0\%\downarrow}$ $\mathop{96.1^{\star}_{\pm 1.4}}\limits_{5.6\%\uparrow}$
w/o Image $\mathop{94.3^{\star}_{\pm 0.5}}\limits_{1.0\%\downarrow}$ $\mathop{90.5^{\star}_{\pm 1.0}}\limits_{1.9\%\downarrow}$ $\mathop{95.1^{\dagger}_{\pm 1.2}}\limits_{1.4\%\uparrow}$ $\mathop{86.5^{\star}_{\pm 2.0}}\limits_{4.9\%\downarrow}$
w/o Video $\mathop{95.0^{\dagger}_{\pm 0.7}}\limits_{0.2\%\downarrow}$ $\mathop{92.1^{\dagger}_{\pm 0.8}}\limits_{0.3\%\downarrow}$ $\mathop{93.0^{\dagger}_{\pm 1.9}}\limits_{0.8\%\downarrow}$ $\mathop{91.1^{\dagger}_{\pm 1.0}}\limits_{0.2\%\uparrow}$
w/o Interaction $\mathop{81.6^{\star}_{\pm 4.5}}\limits_{14.2\%\downarrow}$ $\mathop{77.3^{\star}_{\pm 4.1}}\limits_{16.2\%\downarrow}$ $\mathop{64.4^{\star}_{\pm 5.9}}\limits_{31.3\%\downarrow}$ $\mathop{97.3^{\star}_{\pm 1.2}}\limits_{7.0\%\uparrow}$
w/o Vision $\mathop{94.1^{\star}_{\pm 0.5}}\limits_{1.1\%\downarrow}$ $\mathop{90.3^{\star}_{\pm 1.0}}\limits_{2.2\%\downarrow}$ $\mathop{94.2^{\dagger}_{\pm 1.2}}\limits_{0.4\%\uparrow}$ $\mathop{86.8^{\star}_{\pm 2.1}}\limits_{4.6\%\downarrow}$
w/o Extra $\mathop{78.5^{\star}_{\pm 5.2}}\limits_{17.5\%\downarrow}$ $\mathop{74.5^{\star}_{\pm 4.3}}\limits_{19.3\%\downarrow}$ $\mathop{60.6^{\star}_{\pm 6.0}}\limits_{35.4\%\downarrow}$ $\mathop{97.3^{\star}_{\pm 1.0}}\limits_{6.9\%\uparrow}$
@@ B.5 Expert Performance
Table 7: The performance and temperature of the social bot detectors. The $\star$ indicates that we employ this expert in the final automatic annotator, and -- indicates that temperature scaling is not suitable for this expert. The “All Expert” denotes the ensemble of all experts. The “Annotator (0.75)” denotes that we consider an account a social bot if the likelihood is greater than 0.75.
@@ C.7 Sentiment Variation
We introduce the variation measure to calculate the degree or extent to which public sentiment changes over the news spread. Given a specific information instance and its relation comment, we first calculate the function of the proportion of comments with neutral sentiment over time $f(x)$. We then determine the time series $[x_{0},x_{1},\dots,x_{n}]$, where we set the interval as one hour. The variation is calculated as:
@@ C.9 Semantic Similarity
C.10 Social Bot Comment
We further explore the semantic information of the published content of social bots. Figure 18 illustrates the stances and sentiments of comments published by social bots. For sentiments, social bots would publish far more emotional content than neutral ones. They tend to publish happy and angry content. Namely, the large number of emotional content posted by social bots might affect public sentiment. For stances, bots would publish content with clear stances. Thus we could conclude that social bots would manipulate public opinions by publishing similar and polarized content.

### v3

@@ 4 Misinformation and Social Bots
Social bots are deeply involved in information spread.
We first check the bot percentage:
@@
Misinformation with the same topics has similar content, providing the basis of echo chambers, and social bots amplify this phenomenon.
@@
We first group all pieces of misinformation into clusters with the same topics according to the judgment, where we provide the clustering algorithm in Appendix C.1. We group 11,393 pieces of misinformation into 2,270 clusters, each of which represents a specific topic or event, e.g., “The last two minutes of the air crash”. We aim to explore the textual content similarity of misinformation with the same topics and across different topics.
@@
We first select the 10 largest clusters as representatives, since there is a long-tail effect in cluster size, where we present the selected clusters in Appendix C.2. We visualize the misinformation content representations in Figure 4, which shows the BERT representation using t-SNE dimensionality reduction. It illustrates that the clusters are cohesive, where the silhouette score is 0.29. Namely, each cluster shares similar content while different clusters share significant differences. It suggests that the misinformation environment is homogeneous, providing the basis for echo chambers.
@@
Figure 4: Visualization of misinformation content representations within the largest 10 clusters. Each dot corresponds to a misinformation instance colored according to its topic. The topic labels annotated by the judgment are plotted at each cluster center. We also calculate the silhouette score ($\times 100$). The cohesive clusters indicate misinformation about the same topic having similar content, providing the basis of echo chambers.
@@
We conduct further quantitative analysis by calculating semantic-level and token-level pairwise scores between two instances, where higher scores mean the content of the two instances is more similar. For semantic level, we employ the cosine similarity of the BERT representations, while for token level, we leverage the ROUGE-L score, where we provide the detailed calculation in Appendix C.3. For semantic level, the average value within the same cluster is 0.9448, and the others’ average is 0.5847. For token level, the average value within the same cluster is 0.7815, and the others’ average is 0.0773. We also present completed values in Figure 17 in Appendix C.4. The quantitative results emphasize that misinformation with the same topics has similar content, and misinformation with different topics has distinct content.
@@
We finally explore the patterns of social bots in misinformation. We consider an account a potential echo chamber member if it participates in at least two misinformation discussions (repost, comment, or like) in the same cluster. Figure 5 presents the distribution of bot percentage among echo chamber members and non-members within various clusters. It illustrates that around 18% non-members are social bots. Meanwhile, the members do not contain bots in about half of the clusters. However, in the other half, members exhibit a higher bot percentage across most clusters compared to non-members, reaching up to 50% in many clusters. We speculate that social bots engage in discussions involving misinformation on specific topics, thereby reinforcing the echo chamber effect.
@@
Social bots generate similar content, aiming to manipulate public opinions.
Online information consumers are reluctant to process information deliberately (Möller et al., 2020), becoming susceptible to cognitive biases (Pennycook et al., 2018; Vosoughi et al., 2018). We aim to explore how public opinion changes and how social bots potentially manipulate it. We focus on how public sentiments and stances change in MisBot. We employ two existing classifiers to obtain the sentiments and stances since it is not our contribution, where we provide the details in Appendix C.5.
  • For public sentiments, we categorize sentiments into neutral and non-neutral (including happy, angry, surprised, sad, and fearful). Figure 18 in Appendix C.6 presents sentiment distribution in different social texts. It illustrates that misinformation would publish more emotional content while real information would naturally be reported. On the other hand, public reactions are always emotional, where misinformation shows more anger while real information shows more happiness. Thus, public sentiments are emotional. We further explore the degree or extent to which public sentiments change over the information spread, introducing a variation measure:
@@
    where $f(x_{k})$ denotes neutral sentiment proportion at time $x_{k}$ and we provide the details of $x_{k}$ in Appendix C.7. Figure 6 visualizes sentiment variation distribution, where a larger value means a more drastic change. The average values of misinformation and real information reach 0.287 and 0.225. It illustrates that public sentiment changes are dramatic during information spread.
@@
    Figure 7: The proportion of comments with different stances as the comments increases. Public stances become increasingly polarized, where misinformation contains more comments with clear stances.
@@
Therefore, we can conclude that as the information spreads, public opinions, including sentiment and stance, become polarized, especially regarding misinformation. We then quantitatively prove the correlation between polarization and social bots by the Pearson correlation coefficient: the number of social bots demonstrates strong correlations with the number of comments with non-neutral stances (r = 0.6661) and sentiments (r = 0.6750). We also provide the completed coefficient in Appendix C.8. The relatively high correlation coefficients indicate that social bots might influence public opinion.
Figure 8: Distribution of social bots’ and humans’ semantic similarities, where social bots present higher similarities. Namely, social bots would publish similar content to manipulate public opinion.
We further explore social bot characteristics in information spread. We first calculate the semantic similarity of a specific account, where a higher value means that this account would publish more similar content. We present the detailed calculation method in Appendix C.9 and present the results in Figure 8. It illustrates that social bots generally present higher values than humans. Social bots would publish similar content to amplify the bandwagon effect, where online users adopt behaviors or actions simply because others are doing so, influencing the information spread (Wang and Zhu, 2019). We then identify the sentiments and stances of social texts generated by social bots and present the results in Figure 9. It demonstrates that social bots publish more emotional content and comments with clear stances. The results enhance the finding that social bots generate similar content, aiming to manipulate public opinions.
@@ 5.1 Misinformation Detection
Mainstream detectors focuses on the information content, including text (Hartl and Kruschwitz, 2022; Xiao et al., 2024; Gong et al., 2024), images (Liu et al., 2023a; Zhang et al., 2024b, d), and videos (Tan et al., 2023a; Bu et al., 2024; Zeng et al., 2025). They extract features such as emotion (Zhang et al., 2021; Wan et al., 2025c) and employ neural networks such as graph neural networks (Tao et al., 2024; Zhang et al., 2024f; Lu et al., 2024; Wan et al., 2024b) or neurosymbolic reasoning (Dong et al., 2024) to characterize information. Besides information content, the context, such as user interactions (Shu et al., 2019; Lu and Li, 2020), user profile (Sun et al., 2023; Xu et al., 2024), and evidence (Chen et al., 2024; Wan et al., 2025a) provides helpful signals to detect misinformation. They would model propagation patterns (Cui and Jia, 2024), construct news environments (Yin et al., 2024), or extract multi-hop facts (Zhang et al., 2024a) to enhance detection performance. Recently, to combat LLM-generated misinformation (Zhang et al., 2024e; Venkatraman et al., 2024; Wan et al., 2025b), models employing LLMs (Wan et al., 2024a; Nan et al., 2024) through prompting (Guan et al., 2024; Hu et al., 2024) and in-context learning (Wang et al., 2024) have been proposed.
@@ 6 Conclusion
In this paper, we proposed a novel dataset named MisBot containing information and annotations of misinformation and social bots. MisBot is the most comprehensive; misinformation and real information are distinguishable; and social bots have high annotation quality. Extensive analysis illustrates that (i) social bots are deeply involved in information spread; (ii) misinformation provides the basis of echo chambers, and social bots amplify this phenomenon; (iii) social bots generate similar content aiming to manipulate public opinions.
Acknowledgements
This work is supported in part by the National Natural Science Foundation of China (No. 62192781, No. 62272374), the Natural Science Foundation of Shaanxi Province (No. 2024JC-JCQN-62), the Key Research and Development Project in Shaanxi Province (No. 2023GXLH-024), the Project of China Knowledge Center for Engineering Science and Technology, the Project of Chinese Academy of Engineering “The Online and Offline Mixed Educational Service System for ‘The Belt and Road’ Training in MOOC China”, and the K. C. Wong Education Foundation. This work was completed during the internship at SGIT AI Lab, State Grid Corporation of China.
@@ Acknowledgements
Our proposed dataset is the largest containing misinformation and social bot annotations simultaneously. Meanwhile, it contains multiple modalities, including images and videos, and user interactions. However, due to the focus on news spread, it does not contain interactions like the friend relationship, missing potential relations between social bots and genuine accounts. Meanwhile, we propose a weakly supervised framework for annotating social bots, whose accuracy is comparable to that of crowd-sourcing. However, it struggles to achieve better recall and might miss several social bots. Finally, the experiments in this work focus primarily on the Sina Weibo platform. We expect to expand our analysis to other social media platforms such as $\mathbb{X}$ (Twitter) or Reddit, in future work.
@@
Research on misinformation and social bots is essential for countering online malicious content. This research demonstrates that social bots would amplify the spread of misinformation, enhancing echo chambers and manipulating public opinions. However, it may increase the risk of dual-use, where malicious actors may develop advanced social bots to spread misinformation. We will establish controlled access to ensure that the trained annotator checkpoints are only publicly available to researchers. Meanwhile, we will hide the privacy information in the dataset when we publish it.
@@ A.4 Entity Filter
A.5 Query Method
A.6 Verified Accounts
@@ A.8 Annotation Guideline
Based on the criteria, we write a guideline document for human annotators in Figures 13 and 14. Each human annotator must read this document before annotating.
@@ B.3 Misinformation Detector
where $\mathrm{mean}(\cdot)$ denotes the meaning operator. To encode video, we sample 256 frames from each video and resize each frame into $224\times 224$. We employ pre-trained VideoMAE [^11] Here is the model link. (Tong et al., 2022) $\mathrm{VideoMAE}(\cdot)$ . For each time step, we take 16 frames and set the interval to 12 frames. We could obtain:
@@ B.4 Detector Ablation Study
[…]
@@ B.5 Expert Performance
Table 7: The performance and temperature of the social bot detectors. The $\star$ indicates that we employ this expert in the final automatic annotator, and - indicates that temperature scaling is not suitable for this expert. The “All Expert” denotes the ensemble of all experts. The “Annotator (0.75)” denotes that we consider an account a social bot if the likelihood is greater than 0.75.
@@ C.7 Sentiment Variation
We introduce the variation measure to calculate the degree or extent to which public sentiment changes over the news spread. Given a specific information instance and its relation comment, we first calculate the function of the proportion of comments with neutral sentiment over time $f(x)$. We then determine the time series $x_{0},x_{1},\dots,x_{n}]$, where we set the interval as one hour. The variation is calculated as: